# Yuang $*{}^{*}$thanks: $*{}^{*}$ denotes equal contribution to this paper., Jiawei $*{}^{*}$, Yanqing Zhao, Min Zhang, Mengxin Ren, Xiaofeng Zhao, Xiaosong Qiao, Chang Su, Miaomiao Ma, Hao Yang

### v1

@@ 5 Conclusion
In this paper, we explored combining a speech encoder with an LLM for Chinese ASR and NER tasks using two different architectures: decoder-only and encoder-decoder. Under utterance-level evaluation, both architectures achieved significant improvements over the baseline Conformer model. We further compared these two approaches in terms of their capability to utilize long-form historical information. The long-form evaluations indicate that the decoder-only model benefited more from incorporating history context than the encoder-decoder model. To explain this phenomenon, we conducted a comprehensive analysis of the gate values and attention scores, which revealed the superior ability of the decoder-only model to adjust the importance of speech and text modalities dynamically. For future works, we intend to conduct larger-scale experiments and evaluate our systems on more tasks. We hypothesize that the encoder-decoder model may have an advantage over the decoder-only approach in tasks where audio features are crucial, such as audio event detection. Therefore, we plan to investigate the potential of combining the two approaches to achieve superior performance on a wide range of tasks that require both high-level semantic and fine-grained acoustic information.

### v2

@@ 5 Conclusion
In this paper, we explored combining a speech encoder with an LLM for Chinese ASR and NER tasks using two different architectures: decoder-only and encoder-decoder. Under utterance-level evaluation, both architectures achieved significant improvements over the baseline Conformer model. We further compared these two approaches in terms of their capability to utilize long-form historical information. The long-form evaluations indicate that the decoder-only model benefited more from incorporating historical context than the encoder-decoder model. To explain this phenomenon, we conducted a comprehensive analysis of the gate values and attention scores, which revealed the superior ability of the decoder-only model to adjust the importance of speech and text modalities dynamically. For future works, we intend to conduct larger-scale experiments and evaluate our systems on more tasks. We hypothesize that the encoder-decoder model may have an advantage over the decoder-only approach in tasks where audio features are crucial, such as audio event detection. Therefore, we plan to investigate the potential of combining the two approaches to achieve superior performance on a wide range of tasks that require both high-level semantic and fine-grained acoustic information.